# Hadron-quark hybrid model, modular transformation and Roberge-Weiss transition

### v1

@@ V.2 Calculation of moduli in the fundamental region
As $|\tau|$ increases from 0, the right side of the circle approaches the line $\mathrm{Re}(\tau_{\mathrm{h}})=0.5$, although $x_{0}$ moves the opposite direction. When $T>T_{\mathrm{R}W}$ and $|\tau|>1$, $\tau_{\mathrm{h}}$ forms an arc. Note that Eq. (61) is valid also for $|\tau|>1$ if we replace $\theta$ by $\Delta\phi_{\mathrm{b}-q}$ in Eq. (59). As $|\tau|$ decreases from $\infty$, the arc approaches the line $\mathrm{Re}(\tau_{\mathrm{h}})=0.5$, although $x_{0}$ moves in the opposite direction. This figure also indicates the phase transition at $|\tau|=1$. It is also very interesting that the first equations of (60) and (61) have the same $\theta$-dependence as the Fermi distribution function with imaginary chemical potential, and $\tau=|\tau|e^{i\theta}$ has the same one as the Boltzmann distribution, although the physical meanings of $\tau_{\mathrm{h}}$ and $\tau$ are different from those of the distribution functions. Hence, these equations also indicate that the Boltzmann distribution can be transformed into the Fermi distribution by the modular transformation with the matrix product $\cal{TST}$.
@@ VI Summary
In this paper, we have restricted our discussions on the region where temperature $T$ is above or just below $T_{\rm RW}$. When $T$ is much smaller than $T_{\rm RW}$, the baryon volume approaches the constant value $v_{0}$ and $n_{\rm b}$ is small. Then, the hybrid model reduces to the ordinary hadron resonance gas model Oshima:2023bip.
We also remark that, in this paper, the $\theta$-dependence of the phase $\phi$ of the Polyakov loop $\Phi$ was used as a dynamical input. Usually, the absolute value of $\Phi$ is used to classify the phases of QCD. However, the phase $\phi$ is also a very important quantity. For example, in Refs. Hirakida:2018bkf; Kashiwa:2021ctc, the $\phi$-distribution in the phase space is used to analyze the confinement-deconfinement transition in the framework of persistent homology analyses.

### v4

@@ V.2 Calculation of moduli in the fundamental region
As $|\tau|$ increases from 0, the right side of the circle approaches the line $\mathrm{Re}(\tau_{\mathrm{h}})=0.5$, although $x_{0}$ moves the opposite direction. When $T>T_{\mathrm{R}W}$ and $|\tau|>1$, $\tau_{\mathrm{h}}$ forms an arc. Note that Eq. (63) is valid also for $|\tau|>1$ if we replace $\theta$ by $\Delta\phi_{\mathrm{b}-q}$ in Eq. (61). As $|\tau|$ decreases from $\infty$, the arc approaches the line $\mathrm{Re}(\tau_{\mathrm{h}})=0.5$, although $x_{0}$ moves in the opposite direction. This figure also indicates the phase transition at $|\tau|=1$.
It is also very interesting that the first equations of (62) and (63) have the same $\theta$-dependence as the Fermi distribution function with complex chemical potential, and $\tau=|\tau|e^{i\theta}$ has the same one as the Boltzmann distribution, although the physical meanings of $\tau_{\mathrm{h}}$ and $\tau$ are different from those of the distribution functions. Hence, these equations also indicate that the Boltzmann distribution can be transformed into the Fermi distribution by the modular transformation with the matrix product $\cal{TST}$. When $T\to 0$, the phase transition occurs at $|\tau|=e^{(\mu_{\mathrm{FR}}-m_{\mathrm{F}})/T}=1$ and $p_{\mathrm{F}}=0$ where $m_{\mathrm{F}}$, $\mu_{\mathrm{FR}}$ and $p_{\mathrm{F}}$ are the fermion mass, the real part of the fermion chemical potential, and the Fermi momentum, respectively. That is nothing but the formation of Fermi surface.
@@ VI Summary
In this paper, we have restricted our discussions on the region where temperature $T$ is above or just below $T_{\rm RW}$. When $T$ is much smaller than $T_{\rm RW}$, the baryon volume approaches the constant value $v_{0}$ and $n_{\rm b}$ is small. Then, the hybrid model reduces to the ordinary hadron resonance gas model Oshima et al. (2023).
We also remark that, in this paper, the $\theta$-dependence of the phase $\phi$ of the Polyakov loop $\Phi$ was used as a dynamical input. Usually, the absolute value of $\Phi$ is used to classify the phases of QCD. However, the phase $\phi$ is also a very important quantity. For example, in Refs. Hirakida et al. (2020); Kashiwa et al. (2022), the $\phi$-distribution in the phase space is used to analyze the confinement-deconfinement transition in the framework of persistent homology analyses.
@@ Acknowledgements.
Appendix A Lattice and torus
In this appendix, we brief review the relation between the lattice, the torus and the modular transformation. For a more detailed description, see, eg, the chapter 8 in Ref.  Nakahara (2003).
When $\gamma,\delta\in{\mathbb{C}}-\{0\}$ and $\tau={\delta\over{\gamma}}\notin{\mathbb{R}}$,
$\Lambda=\{m\gamma+n\delta\}~{}(n,m\in{\mathbb{Z}})$ (65)
is called a ”lattice” in the complex plane. (See Fig. 12 )
Fig. 12: The lattice $\Lambda$ in the complex plane. The dots are lattice points.
Equivalently,
$\Lambda_{\tau}=\{m+n\tau~{}(n,m\in{\mathbb{Z}})\},$ (66)
is also a lattice. See Fig.  13.
Fig. 13: The lattice $\Lambda_{\tau}$ in the complex plane. The dots are lattice points.
The lattice $\Lambda^{\prime}$ is called ”similar” to the lattice $\Lambda$ when $\alpha\Lambda^{\prime}=\Lambda$ is satisfied for a nonzero $\alpha\in{\mathbb{C}}$. It is easily shown that $\Lambda_{\tau}$ is similar to $\Lambda$. For the modular transformation (14), it can be also shown that
$\Lambda_{\tau^{\prime}}=(c\tau+d)^{-1}\Lambda_{\tau}.$ (67)
Hence, $\Lambda_{\tau^{\prime}}$ is similar to $\Lambda_{\tau}$.
By identifying the opposite edges of the unit lattice, we can construct the torus $T^{2}$ the complex structure of which is characterized by $\tau$. Then, $\tau$ is called a ”modulus” of the torus. The torus is the quotient space $\mathbb{C}/\Lambda$. When the lattice $\Lambda^{\prime}$ is similar to the lattice $\Lambda$, the complex structure of the torus ${T^{2}}^{\prime}$ is the same as that of $T^{2}$. Hence, the complex structure of the torus is invariant under the modular transformation. By considering the torus, we can extract the topological property which is invariant under the modular transformation.